Calculate the average speed (in m/s) of a car that travels 50km in 30 minutes correct to 1 decimal place.
1
Expert's answer
2020-08-19T12:56:48-0400

The average speed of a car

"v_{\rm ave}=\frac{d}{t}=\frac{50\times 1000\:\rm m}{30\times 60\:\rm s}=27.8\:\rm m/s"